Let R be the relation on X = {1,2,3} defined by (x,y) ∈ R if x<y


  1. List the elements of R (ordered pair)
  2. Find the range of R
  3. Find the domain of R
  4. Draw the Diagraph of R below
  5. Identify the properties of R (kind of graph)
1
Expert's answer
2022-03-18T07:18:18-0400

"1: \left( 1,2 \right) ,\left( 1,3 \right) ,\left( 2,3 \right) \\2:range\left( R \right) =\left\{ 2,3 \right\} \\3:dom\left( R \right) =\left\{ 1,2 \right\}"

4:



5: reflexive - no, symmetric - no, transitive - yes